OpenStudy (anonymous):

Evaluate the expression (4x^3) for x=2

OpenStudy (anonymous):

32

OpenStudy (anonymous):

4(2^3)=4(8)=32

OpenStudy (anonymous):

just plug the 2

OpenStudy (anonymous):

confused got 1024

OpenStudy (anonymous):

Is it \((4x)^3\) or \(4x^3\) ???

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

So first tell me what is the 2*2*2 = ???

OpenStudy (anonymous):

8

OpenStudy (anonymous):

So, the answer is : 4*8 = ???

OpenStudy (anonymous):

\[\huge 4x^3 = 4(2)^3 = 4 \times 8 ?\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome..

OpenStudy (lgbasallote):

hmmm

OpenStudy (anonymous):

hmmmm

OpenStudy (lgbasallote):

careful there...

OpenStudy (anonymous):

Where??

OpenStudy (lgbasallote):

lol nothing. it's right.. just wanted to try that scary catchphrase of @satellite73 ...it always chills my bones when he says that haha

OpenStudy (anonymous):

Ha ha ha... Satellite is just awesome.. When he comes I think all my replies are gonna be wrong..